Run Coordinator report:2025run w3

From GlueXWiki
Jump to: navigation, search

This was a week of general commissioning focussing on trigger development and the commissioning and calibration the FCAL and ECAL.

We started the week with the DAQ having live time issues. There were difficulties decoding the output of the Helicity Decoder and it was added and removed from the DAQ configuration multiple times. The reconstruction code was made more robust to errors. These were solved by the end of the week.

New trigger configurations were trialed. The timing for triggers and detectors was unstable. New timing was determined for most detectors. It was realized that DAQ configurations that accumulates Compton events for the ECAL will not produce sufficient pi0 events for the FCAL. We ran Compton configuration at 100 nA and the pi0 configuration at 300 nA.

Pi0s were observed in the ECAL and FCAL. The first HV determinations for the ECAL using Compton events were implemented. Further Compton iterations took place during the week. The first set voltages for gain balancing the FCAL based on pi0 mass were implemented. We did many other commissioning tasks including a CDC HV scan, TOF HV scan, updating TAGM bias voltages, discriminator thresholds for TAGM and SC, fix to the TOF signal cabling.

There were occasional beam excursions on the active collimator after a beam trip which lasting about 5 minutes. This was mitigated by changes to the beam locks.

We did a partial alignment of diamond JD80-212 which showed a good enhancement peak.